Can you please help me with this linear equation?
An administrator of a popular website is told that a new faster server, can handle 45,000 "hit" (users accessing the site)per second. The web site currently experience a peak demand of about 110,000 hits per second. But, every month, the peak demand increases by 2500 hits per second. Use proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 24 months.

Answer by RAY100(1637)

good problem
increasing demand is an arithmetic sequence
H=110,000+(n-1)*2500
H(24)= 110,000+23*2500=110,000+57500=167,500
proportion of required machines to fit capacity
1/45,000=x/167,500
cross multiply
167,500=45,000x
x=3.7 or 4 machines